The perimeter of a rectangle is 81 cm and the area is 407 cm^2. Set up a quadratic equation and find the dimensions of the rectangle. (show your work)
Answer by checkley79(3341)

81=2L+2W
407=LW OR L=407/W
81=2(407/W)+2W
81=814/W+2W
81=(814+2W*W)/W
81=(814+2W^2)/W
81W=814+2W^2
2W^2-81W+814=0
(2W-37)(W-22)=0
2W-37=0
2W=37
W=18.5 ans.
W-22=0
W=22 ans.
L=407/18.5
L=22 ANS.
L=407/22=18.5 ANS.
PROOF:
81=2*22+2*18.5
81=44+37
81=81